# DROP: Poison Dilution via Knowledge Distillation for Federated Learning

### v1

@@ Abstract.
Federated Learning Security, Targeted Backdoor Attacks, Poisoning Defenses, Knowledge Distillation
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06copyright: nonefootnotetext: Equal Contributionfootnotetext: Correspondence to syros.g@northeastern.edu
1. Introduction
@@
Machine learning (ML) systems greatly benefit from scaling the amount of data used for training ML models (Covert et al., 2024; Parthasarathi et al., 2019; Li et al., 2023; Wu et al., 2021). While centralized training is ideal, data contributors often hesitate to share their raw data with a centralized party. Federated learning (FL) (Kairouz et al., 2021) provides a middle-ground to help alleviate such privacy concerns, while potentially offloading compute onto clients. This design supposedly improves privacy, but it comes at the cost of less control over data and gradient ingestion, allowing malicious clients to introduce malicious goals into the global model through data poisoning (Gu et al., 2019; Wang et al., 2024) or gradient manipulation (Bagdasaryan et al., 2018; Goldblum et al., 2020).
@@
This lack of control over client contributions enables a particularly insidious form of adversarial behavior known as backdoor poisoning attacks (Gu et al., 2019; Liu et al., 2018; Alam et al., 2023; Zhang et al., 2024; Chai et al., 2023). Unlike traditional data poisoning, which aims to degrade overall model performance and mount an availability attack (Rosenfeld et al., 2020; Biggio et al., 2013; Fang et al., 2020; Tolpegin et al., 2020), backdoor attacks embed hidden triggers in the model that lead to selective misclassifications for certain inputs. This allows the global model to perform well on clean data while remaining vulnerable to targeted exploitation. The decentralized nature of FL amplifies this threat, as the server has limited visibility into local training data and model updates, making it difficult to distinguish between honest contributions and adversarial modifications. As a result, backdoor poisoning has emerged as one of the most concerning attack vectors in federated learning.
@@
Although numerous defenses have been proposed to counter poisoning attempts in FL (Yin et al., 2018; Blanchard et al., 2017; Cao et al., 2021; Fung et al., 2018; Shen et al., 2016; Wang et al., 2022; Nguyen et al., 2022; Zhang et al., 2023; Krauß and Dmitrienko, 2023), their evaluation frequently lacks consistency across different federation configurations and adversarial strategies. Many existing works focus on narrow or suboptimal configurations and attack scenarios, limiting their practical applicability (Khan et al., 2023). We identify multiple critical learning configurations—comprising variations in the local number of training epochs, learning rate, and batch size—where existing FL defenses fail to consistently mitigate targeted backdoor attacks. Small changes to these configurations can significantly influence the success of both the attack and the defense, underscoring the need for more robust and generalizable defensive solutions against backdoor poisoning attacks.
@@
In this paper, we address the challenge of defending against targeted backdoor poisoning attacks, a more sophisticated variant of traditional backdoor attacks where the adversary targets a specific “victim” class within the data distribution. This focused strategy makes the attack more stealthy and difficult to detect, as the backdoor trigger affects only a subset of inputs rather than the entire input space. To counter this threat, we propose DROP11 Code available at https://github.com/gsiros/drop (Distillation-based Reduction of Poisoning Signals), a comprehensive, multi-layered defense framework. DROP integrates three key components: (1) Agglomerative Clustering to identify and separate anomalous updates, (2) Activity Monitoring to track and penalize suspicious clients over multiple rounds, and (3) Knowledge Distillation to cleanse the global model using a synthetic dataset guided by the consensus logits from benign client updates. Moreover, we propose DROPlet, a lightweight variant of DROP designed for seamless integration into existing FL defense frameworks. We evaluate DROP against a range of existing defenses across diverse FL configurations and attack scenarios, demonstrating its robustness and adaptability. Unlike existing defenses, which often rely on fixed client-selection strategies or specific attack assumptions, DROP generalizes to dynamic, real-world FL environments where adversarial strategies and learning configurations may vary significantly depending on the learning task. Our results show that DROP effectively mitigates stealthy targeted backdoor attacks, reducing both attack success rates and model degradation across a broad spectrum of configurations.
@@
  • We evaluate seven existing defenses against targeted backdoor attacks and find that most defenses fail to provide adequate protection across a wide range of different learning configurations (learning rate, batch size, number of epochs), exhibiting high sensitivity to changes in the overall FL configuration (Section 3).
  • Motivated by the lack of robust defenses against stealthy targeted backdoors, we propose DROP (Poison Dilution via Knowledge Distillation for Federated Learning), a novel defense mechanism that leverages clustering and activity monitoring techniques with knowledge distillation to tackle adversaries trying to exploit vulnerable learning configurations and different attack strategies (Section 4).
  • Through extensive experimental evaluation across multiple datasets and FL setups, we demonstrate that DROP offers enhanced protection against backdoor attacks from adversaries under varying degrees of stealthiness (Section 5). DROP shows resilience across a wide range of learning configurations and provides consistent attack mitigation throughout training rounds, outperforming existing defenses. For example, we observe an average attack success rate of just $1.93\%$ across 10 diverse FL configurations. However, defending against stealthy attacks under non-IID data distributions remains challenging, with most defenses struggling to offer complete protection in complex dataset scenarios.
Finally, we emphasize the need for researchers to assess robustness across diverse FL configurations (Section 6). We also release our source code as a comprehensive framework that includes implementations of existing defenses, enabling unified comparisons of attack and defense efficacy and facilitating the integration of new attacks and defenses.
2. Background and Threat Model
This section introduces the fundamentals of federated learning, describes the threat of poisoning attacks with a focus on backdoor attacks, and defines the threat model considered in this work.
2.1. Background
2.1.1. Federated Learning
Federated Learning (FL) is a distributed machine learning paradigm where clients collaboratively train a global model while preserving the privacy of their local data, addressing regulatory and privacy concerns (McMahan et al., 2017). The most widely used FL algorithm is Federated Averaging (FedAvg) (McMahan et al., 2017), which operates iteratively through a series of communication rounds.
@@ 2.1.1. Federated Learning
(1) $\mathbf{w}_{t+1}=\frac{1}{|\mathcal{C}_{t}|}\sum_{\begin{subarray}{c}c\in\mathcal{C}_{t}\end{subarray}}\mathbf{w}_{c}^{t}.$
@@ 2.1.2. Poisoning Attacks
Data poisoning, a common attack type, involves adversaries manipulating local training data to influence the global model. This includes injecting mislabeled or out-of-distribution examples to degrade performance (untargeted attacks (Biggio et al., 2013; Jagielski et al., 2018; Mei and Zhu, 2015; Xiao et al., 2015; Shejwalkar and Houmansadr, 2021; Fang et al., 2020)) or subverting predictions for specific tasks (targeted attacks (Koh and Liang, 2017; Shafahi et al., 2018; Suciu et al., 2018)). A more covert form, the backdoor attack (Gu et al., 2019; Bagdasaryan et al., 2018; Wang et al., 2020; Sun et al., 2019), implants hidden behaviors, causing the model to misclassify inputs with a predefined trigger while maintaining accuracy on legitimate data.
Figure 1. Examples of CIFAR-10 images from the plane class with an added backdoor trigger (top-left corner). The presence of the trigger causes the model to misclassify these inputs as the horse class, illustrating the effect of a targeted backdoor attack.
In this work, we focus on targeted backdoor attacks (Chen et al., 2017), where adversaries poison a specific subpopulation (victim class) with a hidden trigger. The backdoor activates only for victim class inputs containing the trigger, remaining covert while exploiting client update variability and the decentralized nature of FL systems, making detection challenging. An example is illustrated in Figure 1.
2.2. Threat Model
@@
Adversary’s Capabilities. The adversary is successful in compromising and gaining control of a fraction $\rho$ out of all clients $\mathcal{N}$ in the system, which is referred to as the Malicious Client Ratio (MCR). This set of compromised clients $\mathcal{N}_{\text{adv}}$ is fixed throughout the federation. Once a client $c\in\mathcal{N}_{\text{adv}}$ is compromised, the adversary gains full control over the client’s local training process, including its dataset $\mathcal{D}_{c}$, training procedure, and model updates $\mathbf{w}_{c}^{t}$ submitted to the server at each round $t$.
@@
(4) $\mathcal{D}_{\text{poisoned}}=\{(\mathbf{x}_{i}+\mathbf{t},y_{\text{target}})\mid(\mathbf{x}_{i},y_{i})\in\mathcal{D}_{\text{victim}}\},$
where $\mathcal{D}_{\text{victim}}$ denotes the set of samples from the victim class. The proportion of poisoned samples relative to the total number of clean samples is referred to as the Data Poisoning Rate (DPR). Additionally, the adversary may alter the client’s update $\mathbf{w}_{c}^{t}$ before submission to the server, embedding adversarial changes that facilitate the backdoor’s propagation to the global model. For convenience, definitions of all acronyms used throughout the paper can be found in Table 1.
@@
Table 1. Glossary of acronyms used throughout the paper.
3. Defenses Falter Against Targeted Backdoors
Several works have proposed defenses against poisoning attacks in FL (Zhang et al., 2023; Nguyen et al., 2022; Blanchard et al., 2017; Yin et al., 2018; Fung et al., 2018; Cao et al., 2021; Wang et al., 2022; Pillutla et al., 2022; Shejwalkar and Houmansadr, 2021). It is standard for such works to focus on a specific learning configuration, for instance setting a specific learning rate and batch size for clients. While demonstrating superior performance in a particular setup is useful, it provides no guarantees about how well the defense would work in other valid learning configurations. We begin with an exploration of some of these learning configurations (Section 3.1) and observe that there exist several equally-valid learning configurations where the model achieves acceptable MTA and the adversary’s objective is preserved, thus making them all equally valid FL configurations. However, we find that all existing defenses we evaluated are effective only for a subset of these valid configurations and no defense is resilient across all configurations (Section 3.2).
3.1. Learning Configuration Matters for Attack Success
@@
Figure 2. Visualizing the impact of the learning configuration (learning rate, batch size, and number of local epochs) on ASR, for 1.25% DPR and 20% MCR, for CIFAR-10 with IID data. We only visualize configurations with MTA $\geq 80\%$. The attack is successful on multiple configurations (in yellow).
3.1.1. Impact of learning rate
@@
3.1.2. Impact of batch size
@@
3.1.3. Impact of training epochs
@@
Table 2 highlights a range of configurations that we term the ”danger zones”—settings where the FL system achieves both high MTA (above 80%) and adequately high ASR (above 85%). These configurations are of great interest because they strike a balance between utility and vulnerability; accurate for legitimate tasks while remaining highly susceptible to targeted backdoor attacks. This highlights the critical need for a defense to be resilient across multiple learning configurations, ensuring robustness regardless of the setup and achieving true learning configuration independence. This is particularly desirable because relying on a single configuration, even if effective, is not always feasible in practice. Various constraints, such as the batch size supported by a specific device, the number of epochs a client is willing to commit to, or other resource limitations, can dictate configurations in practice. Therefore, a defense mechanism that can adapt to different setups without compromising its efficacy is essential for practical and widespread deployment.
3.2. Limitations of Existing Defenses
Based on our analysis, we identify 10 learning configurations where the attack is stealthy (minimal impact on MTA) and highly successful (has ASR higher than 85%), as given in Table 2. For our evaluation we consider key, prominent defense methods which offer diverse strategies to combat backdoor attacks in FL. These defense strategies include coordinate-based approaches like Median (Yin et al., 2018) and Multi-Krum (Blanchard et al., 2017), trust-based methods such as FLTrust (Cao et al., 2021), reputation-based schemes like FoolsGold (Fung et al., 2018) and FLARE (Wang et al., 2022), anomaly detection frameworks like FLAME (Nguyen et al., 2022) and local adversarial training methods like FLIP (Zhang et al., 2023). For more details on these defenses and other related works, see Section 5 and Appendix B.
@@ 3.2. Limitations of Existing Defenses
Table 2. Client FL configurations for successful stealthy attacks on CIFAR-10 i.e., cases with MTA $\geq 80\%$ and ASR $\geq 85\%$.
Evaluation setups for these defenses are often inconsistent, with significant variation in client learning configurations such as learning rate, batch size, and the number of local training epochs. For instance, some works (e.g., FLTrust) specify fixed learning rates and batch sizes, while others (e.g., Multi-Krum, FLAME) provide little to no details about the client training setup, focusing instead on aggregation logic. This lack of standardization in evaluation protocols raises concerns about the reproducibility and generalizability of reported results. A detailed comparison of these evaluation setups is provided in Section C.1.
Figure 3. ASR (%) for our defense (DROP) and various existing defenses for 10 FL configurations (1.25% DPR, 20% MCR) where stealthy attacks are possible. No existing defense provides consistent protection across all configurations.
We find that these approaches often struggle against targeted backdoor attacks due to the attack’s stealthy nature (Figure 3). Median and Multi-Krum fail to exclude the adversary’s updates since the poisoned gradients remain close to the statistical norm, while FLTrust struggles to identify the malicious contributions due to the minimal deviation from expected update patterns. Similarly, FoolsGold proves largely ineffective because the adversary’s contributions exhibit insufficient variability across rounds, making them difficult to detect. FLAME performs reasonably well in certain configurations but fails against attacks that exploit low learning rates. Among the defensive baselines, FLIP is the only one that slightly reduces the ASR, though its mitigation is insufficient to provide robust protection. Existing defenses thus have a severe limitation when it comes to stealthy backdoors: their resilience is sensitive to the specific FL learning configuration.
4. Our Defense: DROP
Figure 4. Overview of the proposed DROP defense. Each round $t$ begins with the server broadcasting the global model to all clients and selecting a subset for local training, which may include both benign (green) and malicious (red) clients. After updates are submitted, DROP employs: (1) Agglomerative Clustering to detect anomalous updates, (2) Activity Monitoring & Penalization to track and penalize suspicious clients, and (3) Knowledge Distillation, where a GAN-generated synthetic dataset and client logits guide the distillation of the global model. The final model $\mathbf{w}_{t+1}$ serves as the global model for round $t+1$.

### v2

@@ Abstract
footnotetext: Equal Contributionfootnotetext: Correspondence to syros.g@northeastern.edu
I Introduction
@@
Machine learning (ML) systems greatly benefit from scaling the amount of data used for training ML models [1, 2, 3, 4]. While centralized training is ideal, data contributors often hesitate to share their raw data with a centralized party. Federated learning (FL) [5] provides a middle-ground to help alleviate such privacy concerns, while potentially offloading compute onto clients. This design supposedly improves privacy, but it comes at the cost of less control over data and gradient ingestion, allowing malicious clients to introduce malicious goals into the global model through data poisoning [6, 7] or gradient manipulation [8, 9].
@@
This lack of control over client contributions enables a particularly insidious form of adversarial behavior known as backdoor poisoning attacks [6, 10, 11, 12, 13]. Unlike traditional data poisoning, which aims to degrade overall model performance and mount an availability attack [14, 15, 16, 17], backdoor attacks embed hidden triggers in the model that lead to selective misclassifications for certain inputs. This allows the global model to perform well on clean data while remaining vulnerable to targeted exploitation. The decentralized nature of FL amplifies this threat, as the server has limited visibility into local training data and model updates, making it difficult to distinguish between honest contributions and adversarial modifications. As a result, backdoor poisoning has emerged as one of the most concerning attack vectors in federated learning.
@@
Although numerous defenses have been proposed to counter poisoning attempts in FL [18, 19, 20, 21, 22, 23, 24, 25, 26], their evaluations lack consistency across different federation configurations and adversarial strategies. Many existing works focus on narrow or suboptimal configurations and attack scenarios, limiting their practical applicability [27]. We identify multiple critical learning configurations—comprising variations in the local number of training epochs, learning rate, and batch size—where existing FL defenses fail to consistently mitigate targeted backdoor attacks. Small changes to these configurations can significantly influence the success of both the attack and the defense, underscoring the need for more robust and generalizable defensive solutions against backdoor poisoning attacks.
@@
In this paper, we address the challenge of defending against targeted backdoor poisoning attacks, a more sophisticated variant of traditional backdoor attacks where the adversary targets a specific “victim” class within the data distribution. This focused strategy makes the attack more stealthy and difficult to detect, as the backdoor trigger affects only a subset of inputs rather than the entire input space. To counter this threat, we propose DROPCode available at https://github.com/gsiros/drop (Distillation-based Reduction of Poisoning Signals), a comprehensive, multi-layered defense framework. DROP introduces a novel architecture that combines effectively three key components: (1) Agglomerative Clustering to identify and separate anomalous updates, (2) Activity Monitoring to track and penalize suspicious clients over multiple rounds, and (3) Knowledge Distillation to cleanse the global model using a synthetic dataset guided by the consensus logits from benign client updates. Through this novel set of countermeasures, DROP provides resilience against both aggressive and stealthy poisoning attacks, across a wide range of learning configurations. Moreover, we propose DROPlet, a lightweight variant of DROP designed for seamless integration into existing FL defense frameworks. We evaluate DROP against seven existing defenses and three poisoning attacks across diverse FL configurations, demonstrating its robustness and adaptability. Unlike existing defenses, which often rely on fixed client-selection strategies or specific attack assumptions, DROP generalizes to dynamic, real-world FL environments where adversarial strategies and learning configurations may vary significantly. Our results show that DROP effectively mitigates stealthy targeted backdoor attacks, reducing attack success rates across a broad spectrum of configurations.
@@
  • We evaluate seven existing defenses against targeted backdoor attacks and find that most defenses fail to provide adequate protection across a wide range of different learning configurations (learning rate, batch size, number of epochs), exhibiting high sensitivity to changes in the overall FL configuration (Section III).
  • Motivated by the lack of robust defenses against stealthy targeted backdoors, we propose DROP (Poison Dilution via Knowledge Distillation for Federated Learning), a novel defense mechanism that leverages clustering and activity monitoring techniques with knowledge distillation to tackle adversaries trying to exploit vulnerable learning configurations and different attack strategies (Section IV).
  • Through extensive experimental evaluation across multiple backdoor attacks, datasets and FL setups, we demonstrate that DROP offers enhanced protection against backdoor attacks from adversaries under varying degrees of stealthiness (Section V). DROP shows resilience across a wide range of learning configurations and provides consistent attack mitigation throughout training rounds, outperforming existing defenses. For example, we observe an average attack success rate of just $1.93\%$ across 10 diverse FL configurations. However, defending against stealthy attacks under non-IID data distributions remains challenging, with most defenses struggling to offer complete protection in complex dataset scenarios.
Finally, we emphasize the need for researchers to assess robustness across diverse FL configurations (Section VI). We also release our source code as a comprehensive framework that includes implementations of existing defenses, enabling unified comparisons of attack and defense efficacy and facilitating the integration of new attacks and defenses.
II Background and Threat Model
We introduce the fundamentals of federated learning, describe the threat of poisoning attacks with a focus on backdoor attacks, and define the threat model considered in this work.
II-A Background
II-A1 Federated Learning
Federated Learning (FL) is a distributed machine learning paradigm where clients collaboratively train a global model while preserving the privacy of their local data, addressing regulatory and privacy concerns [28]. The most widely used FL algorithm is Federated Averaging (FedAvg) [28], which operates iteratively through a series of communication rounds.
@@ II-A1 Federated Learning
$\mathbf{w}_{t+1}=\frac{1}{|\mathcal{C}_{t}|}\sum_{\begin{subarray}{c}c\in\mathcal{C}_{t}\end{subarray}}\mathbf{w}_{c}^{t}.$ (1)
@@ II-A2 Poisoning Attacks
Data poisoning, a common attack type, involves adversaries manipulating local training data to influence the global model. This includes injecting mislabeled or out-of-distribution examples to degrade performance (untargeted attacks [15, 29, 30, 31, 32, 16]) or subverting predictions for specific tasks (targeted attacks [33, 34, 35]). A more covert form, the backdoor attack [6, 8, 36, 37], implants hidden behaviors, causing the model to misclassify inputs with a predefined trigger while maintaining accuracy on legitimate data.
Figure 1: Examples of CIFAR-10 images from the plane class with an added backdoor trigger (top-left corner). The presence of the trigger causes the model to misclassify these inputs as the horse class, illustrating the effect of a targeted backdoor attack.
In this work, we focus on targeted backdoor attacks [38], where adversaries poison a specific subpopulation (victim class) with a hidden trigger. The backdoor activates only for victim class inputs containing the trigger, remaining covert while exploiting client update variability and the decentralized nature of FL systems, making detection challenging. An example is illustrated in Figure 1.
II-B Threat Model
@@
Adversary’s Capabilities. The adversary is successful in compromising and gaining control of a fraction $\rho$ of all clients $\mathcal{N}$ in the system, which is referred to as the Malicious Client Ratio (MCR). This set of compromised clients $\mathcal{N}_{\text{adv}}$ is fixed throughout the federation. Once a client $c\in\mathcal{N}_{\text{adv}}$ is compromised, the adversary gains full control over the client’s local training process, including its dataset $\mathcal{D}_{c}$, training procedure, and model updates $\mathbf{w}_{c}^{t}$ submitted to the server at each round $t$.
@@
$\mathcal{D}_{\text{poisoned}}=\{(\mathbf{x}_{i}+\mathbf{t},y_{\text{target}})\mid(\mathbf{x}_{i},y_{i})\in\mathcal{D}_{\text{victim}}\},$ (4)
where $\mathcal{D}_{\text{victim}}$ denotes the set of samples from the victim class. The proportion of poisoned samples relative to the total number of clean samples is referred to as the Data Poisoning Rate (DPR). Additionally, the adversary may alter the client’s update $\mathbf{w}_{c}^{t}$ before submission to the server, embedding adversarial changes that facilitate the backdoor’s propagation to the global model. For convenience, definitions of all acronyms used throughout the paper can be found in Table I.
@@
TABLE I: Glossary of acronyms used throughout the paper.
III Limitation of Prior FL Backdoor Defenses
Several works have proposed defenses against poisoning attacks in FL [25, 24, 19, 18, 21, 20, 23, 39, 32]. These works usually focus on a specific learning configuration, for instance setting a specific learning rate and batch size for clients. While demonstrating superior performance in a particular setup is useful, it provides no guarantees about how well the defense would work in other valid learning configurations, and thus its sensitivity to a change in configurations. We begin with an exploration of some of these learning configurations (Section III-A) and observe that there exist several equally-valid learning configurations where the model achieves acceptable MTA and the adversary’s objective is preserved, thus making them all equally valid FL configurations. However, we find that all existing defenses we evaluated are effective only for a subset of these valid configurations and no defense is resilient across all configurations (Section III-B).
III-A Learning Configuration Matters for Attack Success
@@
Figure 2: Visualizing the impact of the learning configuration (learning rate, batch size, and number of local epochs) on ASR, for 1.25% DPR and 20% MCR, for CIFAR-10 with IID data. We only visualize configurations with MTA $\geq 80\%$. The attack is successful on multiple configurations (in yellow).
III-A1 Impact of learning rate
@@
III-A2 Impact of batch size
@@
III-A3 Impact of training epochs
@@
Table II highlights a range of configurations that we term the “danger zones”—settings where the FL system achieves both high MTA (above 80%) and adequately high ASR (above 85%). These configurations are of great interest because they strike a balance between utility and vulnerability; accurate for legitimate tasks while remaining highly susceptible to targeted backdoor attacks. This highlights the critical need for a defense to be resilient across multiple learning configurations, ensuring robustness regardless of the setup and achieving true learning configuration independence. This is particularly desirable because relying on a single configuration, even if effective, is not always feasible in practice. Various constraints, such as the batch size supported by a specific device, the number of epochs a client is willing to commit to, or other resource limitations, can dictate configurations in practice. Therefore, a defense mechanism that can adapt to different setups without compromising its efficacy is essential for practical and widespread deployment.
III-B Failure Modes of Prior Defenses
Based on our analysis, we identify 10 learning configurations where the attack is stealthy (minimal impact on MTA) and highly successful (has ASR higher than 85%), as given in Table II. For our evaluation we consider key, prominent defense methods which offer diverse strategies to combat backdoor attacks in FL. These defense strategies include coordinate-based approaches like Median [18] and Multi-Krum [19], trust-based methods such as FLTrust [20], reputation-based schemes like FoolsGold [21] and FLARE [23], anomaly detection frameworks like FLAME [24] and local adversarial training methods like FLIP [25]. For more details on these defenses and other related works, see Section V and Section -B.
@@ III-B Failure Modes of Prior Defenses
TABLE II: Client FL configurations for successful stealthy attacks on CIFAR-10 i.e., cases with MTA $\geq 80\%$ and ASR $\geq 85\%$.
Evaluation setups for these defenses are often inconsistent, with significant variation in client learning configurations such as learning rate, batch size, and the number of local training epochs. For instance, some works (e.g., FLTrust) specify fixed learning rates and batch sizes, while others (e.g., Multi-Krum, FLAME) provide little to no details about the client training setup, focusing instead on aggregation logic. This lack of standardization in evaluation protocols raises concerns about the reproducibility and generalizability of reported results. A detailed comparison of these evaluation setups is provided in Section -C1.
Figure 3: ASR (%) for our defense (DROP) and various existing defenses for 10 FL configurations (1.25% DPR, 20% MCR) where stealthy attacks are possible. No existing defense provides consistent protection across all configurations.
We find that these approaches often struggle against targeted backdoor attacks due to the attack’s stealthy nature (Figure 3). Median and Multi-Krum fail to exclude the adversary’s updates since the poisoned gradients remain close to the statistical norm, while FLTrust struggles to identify the malicious contributions due to the minimal deviation from expected update patterns. Similarly, FoolsGold proves largely ineffective because the adversary’s contributions exhibit insufficient variability across rounds, making them difficult to detect. FLAME performs reasonably well in certain configurations but fails against attacks that exploit low learning rates. Among the defensive baselines, FLIP is the only one that slightly reduces the ASR, though its mitigation is insufficient to provide robust protection. Existing defenses thus have a severe limitation when it comes to stealthy backdoors: their resilience is sensitive to the specific FL learning configuration. In contrast, our proposed defense (which we describe next) generalizes to all learning setups and provides consistent protection.
IV Our Defense: DROP
Figure 4: Overview of the proposed DROP defense. Each round $t$ begins with the server broadcasting the global model to all clients and selecting a subset for local training, which may include both benign (green) and malicious (red) clients. After updates are submitted, DROP employs: (1) Agglomerative Clustering to detect anomalous updates, (2) Activity Monitoring & Penalization to track and penalize suspicious clients, and (3) Knowledge Distillation, where a GAN-generated synthetic dataset and client logits guide the distillation of the global model. The final model $\mathbf{w}_{t+1}$ serves as the global model for round $t+1$.